OpenStudy (anonymous):

I am so stuck on systems of equations, please help 2x+4y=36 10y=5 I need to find x and y

OpenStudy (anonymous):

from equation 10y=5 ten y=1/2 substitute by y=1/2 in the first equation to get x

OpenStudy (anonymous):

thank youuu!

OpenStudy (anonymous):

anytime :)